OpenStudy (destinykiara99):

Find the real numbers x and y that make the equation true. -3 + yi = x + 6i

OpenStudy (mariamikayla):

The value of the coefficient of the complex number 6i must be equal with the coefficient of the complex number yi so that means that y=6 The value of the free terms of the equality must be equal so that means x= -3

jhonyy9 (jhonyy9):

sorry but next to -3 you dont missed an x ?

OpenStudy (destinykiara99):

okay so -3+yi=-3+6i

OpenStudy (destinykiara99):

-3+yi=x+6i -3+yi=x+6i yi=6i x=-3 y=6

OpenStudy (mariamikayla):

-3+yi=x+6i x= -3 so ..... -3+yi=-3+6i but y=6 -3+6i=-3+6i